OpenStudy (anonymous):

Will fan and medal and best response What is 4.8 x 108 m/h in kilometers per hour PLEASE explain how to get the answer

OpenStudy (anonymous):

0.0131674 but i am not sure or it might be 518.4

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

you are welcome

OpenStudy (phi):

is it \[ 4.8 \cdot 10^8 \text{ meters/hour} \] and change it to kilometers/hour ?

OpenStudy (anonymous):

yeah

OpenStudy (phi):

The way to do all these "conversion" problems is to focus on the "units" \[ \frac{meters}{hours} \] and think the same way you do in algebra (if you know algebra) we want to get km/hour, and we will get that by multiplying by "something" \[ \frac{meters}{hours} \cdot \frac{km}{meters} \] in algebra, the meters divided by meters will cancel \[ \frac{\cancel{meters}}{hours} \cdot \frac{km}{\cancel{meters}} = \frac{km}{hour}\]

OpenStudy (anonymous):

alright so then we have to convert the hours to km ?

OpenStudy (phi):

what that means is we need the conversion from km to meters, written in the form km/meters say you know 1000 meters = 1 km we want km/meters so divide both sides by 1000 meters to get \[ \frac{1 \ km}{1000 \ meters} \] that is the fraction we use to do your problem \[ 4.8 \cdot 10^8 \frac {meters}{hour}\cdot \frac{1 \ km}{1000 \ meters} \]

OpenStudy (phi):

notice the "meters" cancel. If we move the numbers all together, we have \[ \frac{4.8 \cdot 10^8}{1000} \frac{km}{hour} \] which is the answer after we simplify the numbers. can you do that ?

OpenStudy (anonymous):

\[\frac{ 4.8 }{ 100000 }\] is what the answer would be ?

OpenStudy (anonymous):

because \[10^{8}/1000=1000000\]

OpenStudy (phi):

10^8 means 10 times itself 8 times. 1000 is 10*10*10, so just the 10's part is \[ \frac{10 \cdot 10\cdot 10\cdot 10\cdot 10\cdot 10\cdot 10\cdot 10}{10 \cdot 10\cdot 10} \] notice we can write that as \[ 10\cdot 10\cdot 10\cdot 10\cdot 10 \cdot \frac{10\cdot 10\cdot 10}{10\cdot 10\cdot 10} \] anything divided by itself is 1, so that simplifies to \[ 10\cdot 10\cdot 10\cdot 10\cdot 10 = 10^5 \] it's faster if you remember the rule "subtract exponents" and do this \[ \frac{4.8 \cdot 10^8}{10^3} \frac{km}{hr} = 4.8 \cdot 10^{8-3} \frac{km}{hr} \\ 4.8 \cdot 10^5 \frac{km}{hr} \]

OpenStudy (anonymous):

oh alright thank you so much !!!!!!!!!!!!!!!